# Uniform Set Systems with Uniform Witnesses

### v1

@@ 1.1. Paper Organization
In Section 2 we present the proof of Theorem 1.4. Then in Section 3 we give the construction that proves Theorem 1.5. Finally, in Section 4 we make some concluding remarks.
@@
Acknowledgments. We thank the organizers of the SLMath workshop “Algebraic and Analytic Methods in Combinatorics” (March 2025) attended by the authors where this project was initiated. We would also like to thank Zixiang Xu for comments on an early draft of this note.
@@
Let $k$ and $r\geqslant 2$ be positive integers. If $\mathcal{F}\subset\binom{[n]}{k}$ has size $|\mathcal{F}|\geqslant k!(r-1)^{k}$, then $\mathcal{F}$ contains an $r$-sunflower.
@@ 2.1. Structural Analysis
In this section, we prove that one can capture the intersection properties in $\mathcal{F}$ using set families of constant size. First, observe that we can interpret the uniform witness condition as follows. For any $B\in\binom{[n]}{s}$ and $F,F^{\prime}\in\mathcal{F}$ such that $B_{F}=B$ and $B\subseteq F^{\prime}$, we must have $(F\setminus B)\cap(F^{\prime}\setminus B)\neq\varnothing$. In other words, the two families $\{F\setminus B\mid F\in\mathcal{F},\,B_{F}=B\}$ and $\{F^{\prime}\setminus B\mid F^{\prime}\in\mathcal{F},\,B\subset F^{\prime}\}$ are cross-intersecting.
@@
    If $F\in\mathcal{F}$, $B\in\mathcal{B}_{1}$ and $E\in\mathcal{E}_{B}$ are such that $E\subset F$, then we have $F=E\cup\{x_{B}\}$.
@@
    $\mathcal{U}(B,B^{\prime})=\left\{E\in{[n]\choose d}\,\middle|\,B\cup B^{\prime}\subset E,~x_{B},x_{B^{\prime}}\not\in E\right\}.$
@@
Finally, to prove Item 3, consider $B,B^{\prime}\in\mathcal{B}_{1}$ with $x_{B}\neq x_{B^{\prime}}$. Assume for contradiction that there exists $E\in\mathcal{U}(B,B^{\prime})\cap\mathcal{E}$, then we must have $E=F^{\prime\prime}\setminus\{x_{B^{\prime\prime}}\}$ for some $B^{\prime\prime}\subset F^{\prime\prime}\in\mathcal{F}$. On the other hand, since $B\subseteq F^{\prime\prime}$, by Item 1 we must have $x_{B^{\prime\prime}}=x_{B}$. Similarly, since $B^{\prime}\subseteq F^{\prime\prime}$, we have $x_{B^{\prime}}=x_{B}$. Therefore we have $x_{B}=x_{B^{\prime\prime}}$ contradicting the assumption that $x_{B}\neq x_{B^{\prime}}$. ∎
@@ 2.3. Stability analysis
Note that for each fixed $B^{\prime}\in\mathcal{B}_{1}^{\prime}$, there are at most $\binom{n-2}{s-1}$ possible $B\in\mathcal{B}_{1}(x_{0})$ such that $x_{B^{\prime}}\in B$. Thus, there are at least $\left\lvert\mathcal{B}_{1}(x_{0})\right\rvert-\binom{n-2}{s-1}$ pairs of the form $(B,B^{\prime})\in\mathcal{P}^{\prime}$ for each $B^{\prime}\in\mathcal{B}_{1}^{\prime}$. From Claims 2.8 and 2.10 with $K=0$, we have
@@
3. A Non-star construction for $d/2<s\leqslant d-1$
In this section, we prove Theorem 1.5.
@@ 3. A Non-star construction for $d/2<s\leqslant d-1$
  • $F\in\mathcal{A}_{2}$: $F=\{2\}\cup A$ for some $A\in\binom{[3,n]}{d}$ such that $|A\cap[3,m]|<s$. We set $B_{F}\subset A$ to be any subset of size $s$ containing $A\cap[m+1,n]$. This is well-defined by the assumption that $|A\cap[m+1,n]|\geqslant d-|A\cap[3,m]|\geqslant d-s+1$ and $s\geqslant d-s+1$.
@@
4. Concluding remarks
In Section 3, the set family we constructed is contained in the union of two stars, namely $F\cap\{1,2\}\neq\varnothing$ for all $F\in\mathcal{F}$ defined in the proof of Theorem 1.5. We were unable to find a construction beyond ones being contained in a union of two stars. This leads us to the following question.
Question 4.1.
Let $d/2<s\leqslant d-2$ and $n\geqslant 2(d+1)$. Does there exist an $s$-witness family $\mathcal{F}\subseteq\binom{[n]}{d+1}$ with witnesses $B_{F}\subseteq F$ of size $s$ for each $F\in\mathcal{F}$ such that
  • $\left\lvert\mathcal{F}\right\rvert=\binom{n-1}{d}$,
  • $F\cap\{1,2,3\}\neq\varnothing$ for each $F\in\mathcal{F}$,
  • $B_{F}\cap\{1,2,3\}=\varnothing$ for the witness $B_{F}$ of each $F\in\mathcal{F}$, and
  • $\mathcal{F}$ is not contained in two stars.
For any set family $\mathcal{F}$, we define $\mathcal{A}_{i}=\{F\in\mathcal{F}\mid F\cap\{1,2,3\}=\{i\}\}$, $\mathcal{A}_{ij}=\{F\in\mathcal{F}\mid F\cap\{1,2,3\}=\{i,j\}\}$, and $\mathcal{A}_{123}=\{F\in\mathcal{F}\mid F\cap\{1,2,3\}=\{1,2,3\}\}$ for any $1\leqslant i<j\leqslant 3$. A potential ‘3-star’ example of size $n-1\choose d$} should be such that $\mathcal{A}_{1}\cup\mathcal{A}_{2}\cup\mathcal{A}_{3}$ is a partition of $\binom{[4,n]}{d}$, $A_{12},A_{13},A_{23}$ covers $\binom{[4,n]}{d-1}$ exactly twice, and $A_{123}=\binom{[4,n]}{d-2}$. We believe that either a positive or a negative answer to this toy problem would lead to a better understanding of 1.3.

### v2

@@ 1. Introduction
We will prove Theorem 1.5 by giving an explicit construction that is contained in the union of two star families but not contained in a single star family. Recall that the piercing number $\tau(\mathcal{F})$ is the cardinality of the smallest set $T$ such that $F\cap T\neq\varnothing$ for all $F\in\mathcal{F}$. In particular, the construction given in the proof of Theorem 1.5 has piercing number $2$. Furthermore, we can generalize the construction in Theorem 1.5 and get $s$-witness families of size $\binom{n-1}{d}$ with arbitrary large piercing number.
Theorem 1.6.
Let $d,s,k\geqslant 1$ be integers such that $d/2<s\leqslant d-k+1$ and $n\geqslant\max(k+(k-1)s,d+2k-1)$. There exists an $s$-witness family $\mathcal{F}\subseteq{[n]\choose d+1}$ of size $n-1\choose d$} such that $\tau(\mathcal{F})=k$.
We remark that our construction gives several pairwise non-isomorphic examples (up to a permutation of the ground set) proving Theorem 1.5, but the construction for Theorem 1.6 gives only one such example (up to a permutation of the ground set). Therefore, we will present both the proof of Theorem 1.5 and the proof of Theorem 1.6 in Section 3 for completeness despite the fact that they share many similarities. [^1] In an earlier arXiv version of this paper, Theorem 1.6 was stated as a question in the special case $k=3$. ChatGPT 5.4 Pro managed to successfully answer this question and the construction in Section 3.2 is a generalization of its response (see https://chatgpt.com/share/69b8e49c-6df0-8009-8180-52161f034bc2 for the conversation).
@@ 1.1. Paper Organization
In Section 2 we present the proof of Theorem 1.4. Then in Section 3 we give the construction that proves Theorems 1.5 and 1.6.
@@
Acknowledgments. We thank the organizers of the SLMath workshop “Algebraic and Analytic Methods in Combinatorics” (March 2025) attended by the authors where this project was initiated. We would also like to thank Zixiang Xu for comments on an early draft of this note. Finally, we thank Zach Hunter for pointing out typos in the first draft posted on arXiv.
@@
Let $k$ and $r\geqslant 2$ be positive integers. If $\mathcal{F}\subseteq\binom{[n]}{k}$ has size $|\mathcal{F}|\geqslant k!(r-1)^{k}$, then $\mathcal{F}$ contains an $r$-sunflower.
@@ 2.1. Structural Analysis
In this section, we prove that one can capture the intersection properties in $\mathcal{F}$ using set families of constant size. First, observe that we can interpret the uniform witness condition as follows. For any $B\in\binom{[n]}{s}$ and $F,F^{\prime}\in\mathcal{F}$ such that $B_{F}=B$ and $B\subseteq F^{\prime}$, we must have $(F\setminus B)\cap(F^{\prime}\setminus B)\neq\varnothing$. In other words, the two families $\{F\setminus B\mid F\in\mathcal{F},\,B_{F}=B\}$ and $\{F^{\prime}\setminus B\mid F^{\prime}\in\mathcal{F},\,B\subseteq F^{\prime}\}$ are cross-intersecting.
@@
    If $F\in\mathcal{F}$, $B\in\mathcal{B}_{1}$ and $E\in\mathcal{E}_{B}$ are such that $E\subseteq F$, then we have $F=E\cup\{x_{B}\}$.
@@
    $\mathcal{U}(B,B^{\prime})=\left\{E\in{[n]\choose d}\,\middle|\,B\cup B^{\prime}\subseteq E,~x_{B},x_{B^{\prime}}\not\in E\right\}.$
@@
Finally, to prove Item 3, consider $B,B^{\prime}\in\mathcal{B}_{1}$ with $x_{B}\neq x_{B^{\prime}}$. Assume for contradiction that there exists $E\in\mathcal{U}(B,B^{\prime})\cap\mathcal{E}$, then we must have $E=F^{\prime\prime}\setminus\{x_{B^{\prime\prime}}\}$ for some $B^{\prime\prime}\subseteq F^{\prime\prime}\in\mathcal{F}$. On the other hand, since $B\subseteq F^{\prime\prime}$, by Item 1 we must have $x_{B^{\prime\prime}}=x_{B}$. Similarly, since $B^{\prime}\subseteq F^{\prime\prime}$, we have $x_{B^{\prime}}=x_{B}$. Therefore we have $x_{B}=x_{B^{\prime\prime}}$ contradicting the assumption that $x_{B}\neq x_{B^{\prime}}$. ∎
@@ 2.3. Stability analysis
Note that for each fixed $B^{\prime}\in\mathcal{B}_{1}^{\prime}$, there are at most $\binom{n-2}{s-1}$ possible $B\in\mathcal{B}_{1}(x_{0})$ such that $x_{B^{\prime}}\in B$. Thus, there are at least $\left\lvert\mathcal{B}_{1}(x_{0})\right\rvert-\binom{n-2}{s-1}$ pairs of the form $(B,B^{\prime})\in\mathcal{P}^{\prime}$ for each $B^{\prime}\in\mathcal{B}_{1}^{\prime}$. Recall that we defined $M\geqslant 0$ such that $|\mathcal{B}_{1}(x_{0})|=|\mathcal{B}_{1}|-M$. From Claims 2.8 and 2.10 with $K=0$, we have
@@
3. Non-star constructions
We first prove Theorem 1.5 in Section 3.1, and then prove Theorem 1.6 in Section 3.2
3.1. A non-star construction for $d/2<s\leqslant d-1$
Now we give a construction that proves Theorem 1.5.
@@ 3.1. A non-star construction for $d/2<s\leqslant d-1$
  • $F\in\mathcal{A}_{2}$: $F=\{2\}\cup A$ for some $A\in\binom{[3,n]}{d}$ such that $|A\cap[3,m]|<s$. We set $B_{F}\subseteq A$ to be any subset of size $s$ such that $|B_{F}\cap[m+1,n]|\geqslant d-s+1$. This is well-defined by the assumption that $|A\cap[m+1,n]|\geqslant d-|A\cap[3,m]|\geqslant d-s+1$ and $s\geqslant d-s+1$.
@@
3.2. A non-star construction with arbitrarily large piercing number
The construction for Theorem 1.6 is similar to the one given in Section 3.1. In particular, when the piercing number $k=2$, the following construction recovers the one in Section 3.1 with $m=s+2$.
Proof of Theorem 1.6.
Let $U=[n]\setminus[k]=\{k+1,\ldots,n\}$. Let $S_{1},S_{2},\ldots,S_{k-1}\subseteq U$ be a collection of pairwise disjoint subsets of size $s$. This can be done provided that $n\geqslant k+(k-1)s$. For $i\in[k-1]$ and $m\geqslant s$, let us define the families
$\mathcal{A}_{i}(m)\overset{\rm{def}}{=}\left\{A\in{U\choose m}\,\middle|\,S_{i}\subseteq A\right\},$
and
$\mathcal{A}_{k}(m)\overset{\rm{def}}{=}\left\{A\in{U\choose m}\,\middle|\,S_{i}\not\subseteq A~\forall i=1,\ldots,k-1\right\}.$
Using the assumption that $2s>d$, we know that these families are pairwise disjoint and they cover the sets in $U\choose m$} if $m\leqslant d$. For a non-empty subset $I\subseteq[k]$, we define
$\mathcal{A}_{I}\overset{\rm{def}}{=}\bigsqcup_{i\in I}\mathcal{A}_{i}\left(d+1-|I|\right)\subseteq{U\choose d+1-|I|}$.
Let $\mathcal{F}_{I}\overset{\rm{def}}{=}\{I\cup A\mid A\in\mathcal{A}_{I}\}$ and define $\mathcal{F}\overset{\rm{def}}{=}\bigsqcup_{\varnothing\neq I\subseteq[k]}\mathcal{F}_{I}$. We claim that $\mathcal{F}$ is an $s$-witness family of size $n-1\choose d$} with piercing number $k$. First, let us compute the size of $\mathcal{F}$. Note that for each $t=1,\ldots,k$, we have
$\sum_{I\in{[k]\choose t}}|\mathcal{A}_{I}|=\sum_{i=1}^{k}{k-1\choose t-1}|\mathcal{A}_{i}|={k-1\choose t-1}{n-k\choose d+1-t}$,
and so
$|\mathcal{F}|=\sum_{\varnothing\neq I\subseteq[k]}|\mathcal{A}_{I}|=\sum_{t=1}^{k}{k-1\choose t-1}{n-k\choose d+1-t}={n-1\choose d}$.
Next, we show that $\mathcal{F}$ is indeed an $s$-witness family. We begin by defining the witness sets $B_{F}$ for each $F\in\mathcal{F}$. Suppose that $F\in\mathcal{F}_{I}$ with nonempty $I\subseteq[k]$, so that there exists $A\in\mathcal{A}_{I}$ such that $F=I\cup A$. Then we define $B_{F}$ in the following way:
  • If $S_{i}\subseteq A$ for some $i\in[k-1]$, then we set $B_{F}=S_{i}$.
  • Otherwise, let $J$ be the set of indices $j\in[k-1]\setminus I$ such that $A\cap S_{j}\neq\varnothing$. For every $j\in J$, fix an arbitrary element $x_{j}\in S_{j}\cap A$. Let $B_{F}$ be an arbitrary subset of $A\setminus\{x_{j}\mid j\in J\}$ of size $s$.
Note that in the second case we have $|J|\leqslant k-|I|$. Thus, we have
$|A\setminus\{x_{j}\mid j\in J\}|=|A|-|J|=d+1-|I|-|J|\geqslant d+1-k\geqslant s,$
and hence it is possible to pick an $s$-subset $B_{F}\subseteq A\setminus\{x_{j}\mid j\in J\}$. Furthermore, we note that $B_{F}\subseteq U$ for all $F\in\mathcal{F}$.
Now, we check the $s$-witness condition. That is, we check that $F\cap F^{\prime}\neq B_{F}$ for all $F,F^{\prime}\in\mathcal{F}$. Suppose that $F=I\cup A$ and $F^{\prime}=I^{\prime}\cup A^{\prime}$ with nonempty $I,I^{\prime}\subseteq[k]$, $A\in\mathcal{A}_{I}$, and $A^{\prime}\in\mathcal{A}_{I^{\prime}}$. We check the condition in the following case analysis.
  • If $I\cap I^{\prime}\neq\varnothing$, then $F\cap F^{\prime}\supseteq I\cap I^{\prime}$. Since $I\cap I^{\prime}$ is disjoint from from $B_{F}\subseteq U$, we have $F\cap F^{\prime}\neq B_{F}$ in this case.
  • If $I\cap I^{\prime}=\varnothing$ and $S_{i}\subseteq A$ for some $i\in[k-1]$, then $i\in I$ and $B_{F}=S_{i}$. It follows that $i\not\in I^{\prime}$ and hence $B_{F}=S_{i}\not\subseteq A^{\prime}$.
  • Otherwise, we have $I\cap I^{\prime}=\varnothing$ and $S_{i}\not\subseteq A$ for all $i\in[k-1]$. It follows that $k\in I$. Since $A^{\prime}\in\mathcal{F}_{I^{\prime}}$ and $I^{\prime}\subseteq[k-1]$, there exists $i^{\prime}\in I^{\prime}$ such that $S_{i^{\prime}}\subseteq A^{\prime}$.
    • If $A\cap S_{i^{\prime}}=\varnothing$, then we have
      $|F\cap F^{\prime}|=|A\cap A^{\prime}|=|A\cap(A^{\prime}\setminus S_{i^{\prime}})|\leqslant|A^{\prime}|-|S_{i^{\prime}}|\leqslant d-s<s,$
      and hence $F\cap F^{\prime}\neq B_{F}$.
    • Otherwise, we have $A\cap S_{i^{\prime}}\neq\varnothing$. By construction there is an element $x_{i^{\prime}}\in A\cap S_{i^{\prime}}$ which does not lie in $B_{F}$. On the other hand, since $S_{i^{\prime}}\subseteq A^{\prime}$, we clearly have $x_{i^{\prime}}\in A\cap A^{\prime}\subseteq F\cap F^{\prime}$. Therefore, $F\cap F^{\prime}\neq B_{F}$.
By the case analysis above, we know that $\mathcal{F}$ is an $s$-witness family.
It remains to show that $\tau(\mathcal{F})=k$. It is clear that $[k]$ pierces every set in $\mathcal{F}$, and hence $\tau(\mathcal{F})\leqslant k$. Now, assume for the sake of contradiction that there is some (k1)(k-1)-set $T=I\sqcup R$, where $I\subseteq[k]$ and $R\subseteq U$, that intersects every set in $\mathcal{F}$. For any element $j\in[k]\setminus I$, we know that $R$ intersects every set in $\mathcal{A}_{j}(d)$. For each $j\in[k-1]\setminus I$, then we must have $R\cap S_{j}\neq\varnothing$. Otherwise, we can find a set $A\subseteq U\setminus R$ of size $d$ such that $S_{j}\subseteq A$ and $A\cap R=\varnothing$. This can be done since $\left\lvert U\setminus R\right\rvert=\left\lvert U\right\rvert-\left\lvert R\right\rvert\geqslant n-k-(k-1)\geqslant d$. Therefore, we know that $\left\lvert(I\cup R)\cap S_{j}\right\rvert=1$ for all $j\in[k-1]$, and $k\notin I$. Thus, $R$ must intersect every set in $\mathcal{A}_{k}(d)$. However, we can find a set $A\subseteq U\setminus R$ of size $d$ that does not contain any $S_{i}$ with $i\in[k-1]$ as a subset. This is because we only need to avoid the set $R$ and also avoid one element from $S_{i}$ for each $i\in I$. This is doable since $\left\lvert U\right\rvert-\left\lvert R\right\rvert-\left\lvert I\right\rvert=n-k-(k-1)\geqslant d$.
We can conclude that $\mathcal{F}$ is an $s$-witness family of size $\binom{n-1}{d}$ and it has piercing number $k$. This completes the proof. ∎